Colloidal Clusters: The crew enabled the UV (ultraviolet) illumination system for the Colloidal Clusters sample bags. UV light is used to activate the UV-curable gelling agent and terminate the experiment. Following the appropriate UV exposure time, the crew performed steps needed to close out the experiment. JAXA’s Colloidal Clusters investigates the mechanism for formation of tetrahedral clusters of negatively and positively charged particles suspended in liquid. These clusters are examined and characterized after return to Earth. The clusters may be useful as building blocks of future photonic, or light-manipulating, materials.
Droplet Formation Study (DFS): The crew performed the setup and initiation of recording for additional runs of the DFS investigation. The team is still discussing issues with data archival to Blue-ray disc. It’s believed only a partial set of data was collected for one of the runs, but each testpoint is performed in duplicate so this is likely a minimal impact. Droplet Formation Studies in Microgravity (Droplet Formation Study, or more simply DFS) evaluates water droplet formation and water flow of Delta Faucet’s H2Okinetic shower head technology. To conserve water, flow rates in shower devices have been reduced, but this lower flow rate also reduces the effectiveness of these devices and often causes consumers to take longer showers, undermining the goal of using less water. Gravity’s full effects on formation of water droplet size are unknown, and research in microgravity could help improve the technology, creating better performance and improved user experience while conserving water and energy.
ISS HAM Pass: The crew participated in an ISS HAM pass with the American School of Rio de Janeiro, Rio de Janeiro, Brazil. The school, which began in 1937 in a small house in Ipanema, has since expanded to two amazing world-class campuses. They have been proud to help generations of young people find their own unique path, and today are privileged to be able to offer a once in a lifetime opportunity – a chance to talk to astronauts in space! ISS Ham Radio provides opportunities to engage and educate students, teachers, parents and other members of the community in science, technology, engineering and math by providing a means to communicate between astronauts and the ground HAM radio units.
EHS Benzene Chip Measurement System (CMS) Operations: As part of the continued benzene investigation, the crew took two benzene measurements in the Lab today with the CMS reader. Results were provided to the ground. Readings to date have been below the detectable limits. The crew will continue to take samples in the LAB every other day and the frequency may be altered based on real-time indications.
In Flight Maintenance (IFM) Crew Quarters (CQ) Deck and Overhead Cleaning: A USOS crewmember cleaned the Overhead and Deck CQ today. During this activity, the crew cleaned behind panels, intake and exhaust ducts, fans and airflow sensors as required.
Cabin Fan Assembly 1 (CFA1) Remove and Replace (R&R): Crew continued with the second day of operations to replace the cabin fan assembly in the COL1D1 rack. Today the crew rotated the rack and performed the cabin fan assembly R&R. Prior to tilting the rack, the crew uninstalled the Light Ions Detector (LIDAL). During the rack rotation the crew encountered issues with a fastener that was no longer captive. The ground requested the crew label the damaged components. The crew was able to proceed and complete the scheduled rack rotation activities. ESA confirmed successful activation of the fan.
Completed Task List Activities:
- Data Prep for Return-DM-2 Crew (on-going)
Today’s Ground Activities:
All activities are complete unless otherwise noted.
- Ground Support for MT Translate
- Ground Support for JEM RMS Ops
- RAIU 3 PBIT Clear
Look Ahead Plan
Friday, 7/31 (GMT 213)
- BioMole Part 3
- Droplet Formation Study
- MELFI Icebrick Insert
- Radi-N2 Deploy
- Dragon Prepack Transfer
- OBT Dragon Depart Review
- EML Installation
Saturday, 8/1 (GMT 214)
- BioMole Stow
- GLACIER Transfer
- Emergency Equipment Transfer
- Dragon Cargo Ops/Closeout
- Dragon IDA Egress
- Dragon Undock
Sunday, 8/2 (GMT 215)
- No Utilization Activities
- Crew Off Duty
Today’s Planned Activities:
All activities are complete unless otherwise noted.
- ТБУ. Check ТБУ-В #07 Thermostat Temperature and Keep Out Zones in front of ТБУ-В Inlets and Outlets
- Droplet Formation Study Water Charging, Test Operations and Cleanup
- Light Ions Detector Uninstall & Reinstall
- Colloidal Clusters Ultra Violet Irradiation
- KONSTANTA-2. Session 2 Setup and Ops
- Crew Departure Preparations for Return to Earth
- COL1D1 Rack Rotate Part
- EHS Benzene Chip Measurement System (CMS) Operations
- URAGAN. Observation and Photography using Photo Equipment
- Cabin Fan Assembly 1 R&R
- SCENARIO. Observation and Photography of Catastrophic Events and their Impact
- Columbus Deck Bags restow under COL1D1 Rack.
- Colloidal Clusters Experiment Stop
- ASEPTIC. Surface Samples Collection and Air Samples Collection
- Urine Transfer System Backup EDV Swap
- BIOMAG-M. Transfer and Insert into the Thermostat at +4 С
- Cargo Transfer from DC1 Progress #444 (76P)
- МО-21. MEDIUM 1 Results
- In Flight Maintenance (IFM) Crew Quarters (CQ) Deck & Overhead Cleaning
- Health Maintenance System (HMS) Automated External Defibrillator (AED) Inspection
- Installation of Cable Insert to Power On [B3] Fan in DC1
- Food Physiology Crew Diet Briefing
- ISS HAM Service Module Pass
- Polar Sample Transfer 1 From Cold Stowage
- Glacier Sample Transfer From MELFI
- URAGAN. SOVA Equipment using Photo Equipment